OpenStudy (voltage):

looking for a quick fan and medal help me and you got it. Multiply (−7)(2)(−2). −28 −14 28 14

OpenStudy (anonymous):

It would be 28. You are multiplying them. -7*2 = -14 -14*-2 = 28